### v1

# The NHIM development scenario in an asymmetric binary system of dwarf galaxies

Abstract
We study the development scenario of a three-degree-of-freedom Hamiltonian system. We observe how the fundamental normally hyperbolic invariant manifolds (NHIMs) form a homoclinic/heteroclinic tangle. As the energy increases, the NHIMs begin to lose normal hyperbolicity and decay, creating tangential transient effects. We also observe a certain kind of coordination between the development scenarios of these NHIMs. We analyse this phenomenon using Poincaré maps and phase space indicator functions. As an example of demonstration, we use a relatively simple model of a test particle moving in the gravitational field of a pair of interacting dwarf galaxies.
1 Introduction
In dynamical systems showing chaos, unstable invariant subsets are the geometrical structures in the phase space which determine the global dynamics and, in particular, the creation of chaos to a large extent. As the simplest example, we can think of the unstable fixed points in a 2-dimensional map whose stable and unstable manifolds build up the chaotic invariant set, which can be displayed pictorially as a horseshoe construction traced out by the intersecting stable and unstable manifolds of the most important unstable fixed points. In maps with a domain of higher dimension, the role played in 2-dimensional maps by unstable fixed points is taken over by normally hyperbolic invariant manifolds (standard abbreviation NHIMs) of codimension 2; for general information on NHIMs see [1]. These subsets have stable and unstable manifolds of codimension 1, which can form walls, channels, and tubes in the domain of the map and thereby direct the global dynamics. This is the direct generalisation of the role of stable and unstable manifolds of hyperbolic fixed points in 2-dimensional maps. Then it is easy to understand that the parameter dependence of the central NHIMs has an enormous influence on the parameter dependence of the whole dynamics. Therefore, we are interested in a good understanding of the typical development scenarios of NHIMs under parameter changes. In autonomous Hamiltonian systems (the class of systems we deal with in the present article), the most important parameter is in many cases the total energy $E$, i.e. the conserved numerical value of the generating function $H$ of the dynamics.
Hamiltonian systems have an effective potential; for the general recipe to construct it for a given Hamiltonian function, see [2, 3, 4]. And usually, the most important NHIMs of codimension 2 sit over the index-1 saddle points of this effective potential. Then these NHIMs act as transition states over these saddles and direct the flow over these saddles [5, 6, 7]. We can imagine that from parts of the stable and unstable manifolds of the NHIMs, we glue together tubes (cylinders) which guide the flow over the saddle of the effective potential. This property demonstrates the importance of the NHIMs in the generation of the global properties of the dynamics.
For Hamiltonian systems with three degrees of freedom (3-dof), we are at the moment far away from a complete understanding of the typical development scenarios of codimension-2 NHIMs. We only have hard mathematical persistence theorems which state that NHIMs survive perturbations as long as the normal instability remains larger than the tangential instability, for various versions of proofs see [8], [9], chapter 3 in [1], and chapter 3 in [10]. Many numerical examples indicate that NHIMs really decay in regions where the tangential instability surpasses the normal instability. In the present situation, it is worth studying many examples of development scenarios of NHIMs numerically to gain more experience with the typical and frequent development scenarios.
There are some investigations of the loss of normal hyperbolicity of invariant surfaces [11, 12, 13, 14]. Unfortunately, these publications use 2-dof systems or use a simplification which is equivalent to the reduction to 2 essential degrees of freedom. And they are mainly concerned with the consequences of the loss of normal hyperbolicity to transition state theory and to the flow rates over the saddles. They are not concerned with the disintegration of the NHIM surface into (partially lower-dimensional) invariant fragments, a topic into which we will delve in the present article. For some further observations of qualitative changes of NHIMs under parameter changes, see [15, 16, 17, 18].
NHIMs of dimension 2 or more in a map have an internal dynamics which can be displayed graphically by plotting the Poincaré map restricted to the NHIM itself, which we will call the restricted map for short. For an early example of the use of such a restricted map, see Fig.1 in [11]. It is the same basic idea as the restriction of the Poincaré map to the centre manifold of the saddle point, see for example Fig.6 in [19]. For maps with 4 dimensions in total, this restricted map for NHIMs of codimension 2 is itself 2-dimensional and can be plotted conveniently. This restricted map, as a function of the parameters of the system, is an excellent way to display graphically the development scenario of the NHIMs. The numerical method to follow the NHIMs under parameter changes and to display the restricted map have been explained in detail in [20].
In various examples, we have observed that in systems with several fundamental NHIMs the development scenarios of these NHIMs are coordinated and important qualitative changes of one particular NHIM occur in the same small parameter interval in which also some other NHIMs suffer important changes. One of these examples has been a simple model for the motion of a test particle in the effective potential of a rotating symmetric system of two spherical dwarf galaxies [21]. This system has three collinear Lagrange points and codimension-2 NHIMs over these saddle points of the effective potential. The two outer NHIMs are symmetry-related, whereas the middle one is qualitatively different from the outer ones. Nevertheless, the development scenario of the middle NHIM is coordinated with the development scenario of the two outer NHIMs. Now the reader can ask whether the discrete left-right symmetry of the complete system is responsible for this coordination of the development scenarios. Motivated by this question, we investigate in the present article the asymmetric generalisation of the model from [21]. In the present article, we set the masses of the two spherical dwarf galaxies to different values in contrast to [21], where these two masses have been equal.
Another example of the occurrence of coordination effects in the development scenario of NHIMs has been observed in [22]. This system is the motion of a test particle in a rotating double-barred galaxy. The model has four codimension-2 NHIMs over index-1 saddles of the effective potential. These four NHIMs fall into two nonequivalent groups, where the two NHIMs within each group are symmetry related. Nevertheless, we found a coordination of the scenarios of all four NHIMs, i.e. also between the nonequivalent groups.
In the treatment of these previous systems, we did not observe any important role of the transient outer parts of partially decayed NHIMs. In the meantime, two interesting cases of such transient effects have been observed, see [23] and [24]. Interestingly, in the asymmetric case of the two dwarf galaxies, the topic of the present article, the transient outer parts of the decaying NHIMs influence the coordination of the development and decay scenario and thereby of the creation and development of chaos. This is a new event, not noticed in previous studies. We consider it a remarkable new result of the investigation presented here. For general information on transients and in particular on transient chaos, see the textbook [25].
In section 2, we introduce the asymmetric model and explain the important extremal points of its effective potential. Section 3 displays graphically the development scenario of the 3 NHIMs and of the most important periodic orbits related to these NHIMs. In section 4, we have a closer look at the creation of chaos and at heteroclinic connections between NHIMs, including the role of transient effects. In section 5, we use the delay time as an indicator function to obtain more information about the phase space structure, and in section 6 we draw conclusions and give final remarks.
2 The model
In [21], the functional form of the model has already been given for general values of the masses $M_{1}$ and $M_{2}$ of the two galaxies. Therefore, we give here only a brief repetition of the most important properties of the model. The total mass of the two galaxies taken together is $M_{t}=M_{1}+M_{2}$. The distance between the centres of the two galaxies is $R$. We describe the dynamics in a corotating frame with Cartesian coordinates $x,y,z$ where the origin is the centre of mass of the whole system. The $z$ axis is the axis of rotation with the angular velocity $\Omega$. In the rotating frame, the centres of the two individual galaxies are located along the $x$ axis at the points $x_{1}=-RM_{2}/M_{t}$ and $x_{2}=RM_{1}/M_{t}$ respectively and independently of the time. The gravitational potential for a test particle at a general point $x,y,z$ is given as
$V_{g}=-\frac{GM_{1}}{r_{1}}-\frac{GM_{2}}{r_{2}},$ (1)
where $G$ is the gravitational constant and
$r_{1}^{2}=(x-x_{1})^{2}+y^{2}+z^{2}+c^{2},$ (2)
$r_{2}^{2}=(x-x_{2})^{2}+y^{2}+z^{2}+c^{2}.$
Here $c$ is a size measure for the galaxies, which we take equal for the two galaxies. The two galaxies rotate around each other and when $R$ is large compared to $c$ then we can approximate the rotational angular velocity $\Omega$ by the Kepler value
$\Omega=\sqrt{\frac{GM_{t}}{R^{3}}}.$ (3)
Accordingly, the effective potential in the rotating frame is given as
$V_{eff}(x,y,z)=V_{g}(x,y,z)-\frac{\Omega^{2}}{2}(x^{2}+y^{2}).$ (4)
We use the same units as in the symmetric case; they are explained in detail in section 2 of [21].
In these units, the parameters of the system used in the following are $c=0.25$, $R=6$, $M_{1}=0.06$ and $M_{2}=0.14$. Note that accordingly $M_{t}=0.2$, exactly the same value as the one used for the symmetric case in [21]. Also, the values of $c$ and $R$ are chosen the same as in the symmetric case. We hope that this makes a comparison between the symmetric and asymmetric cases easier. The momenta conjugate to the Cartesian position coordinates are $p_{x},p_{y},p_{z}$. In these coordinates, the generating function of the dynamics (the Hamiltonian function) is given as
$H(x,y,z,p_{x},p_{y},p_{z})=\frac{1}{2}(p_{x}^{2}+p_{y}^{2}+p_{z}^{2})+V_{g}(x,y,z)-\Omega L_{z}.$ (5)
Here $L_{z}=xp_{y}-yp_{x}$ is the $z$ component of the angular momentum. We describe the dynamics in a corotating system. Therefore, the numerical value of the function $H$ is not the energy of the test particle as measured by an observer in an inertial system. This energy, measured by the inertial observer, is not conserved because the potential seen by the inertial observer is explicitly time dependent. For the inertial observer, the conserved generating function from Eq.5 is the Jacobi conserved quantity. In the following, we use the letter $E$ for the numerical value of the function $H$, and call this number the energy, even though it is not the energy as measured by the inertial observer.
In later sections, a knowledge of some symmetries of this system will be useful. First, we notice that the effective potential is symmetric under $y\rightarrow-y$ and under $z\rightarrow-z$. If a trajectory starts with $z=0$ and $p_{z}=0$, then this trajectory will lie completely in the horizontal plane $z=0$ of the position space. Therefore, this horizontal plane is an invariant 2-dof subsystem. Because we describe the dynamics in a rotating frame, we do not find the usual time reversal symmetry. However, if we combine the transformation $t\rightarrow-t$ with a reflection in the plane $y=0$, then the orientation of rotation is also reversed and the Hamiltonian equations of motion remain the same. Therefore, the whole dynamics is invariant under the following transformation
$(t,x,y,z,p_{x},p_{y},p_{z})\rightarrow(-t,x,-y,z,-p_{x},p_{y},-p_{z}).$ (6)
This transformation can be used to transform stable manifolds of invariant subsets into unstable manifolds of the corresponding $y$-reflected invariant subsets. This transformation will become useful in subsection 4b.
Some contour lines of the effective potential from Eq.4 are plotted in the horizontal xyx-y plane in Fig.1. The points $L_{1}$, $L_{2}$, and $L_{3}$ are the collinear Lagrange points. In the 3-dimensional position space, they are index-1 saddles of $V_{eff}$ where the $x$ direction is the unstable direction. The 3 important NHIMs of the system sit over these saddle points. The Lagrange points $L_{4}$ and $L_{5}$ are index-2 saddle points of $V_{eff}$ and are of little importance in the following. The points $P_{1}$ and $P_{2}$ are relative minima of $V_{eff}$ and they lie close to the centres of the two galaxies because the distance $R$ is large compared to the size $c$. Compare this plot with Fig.3 in [21], which displays the corresponding symmetric case. In the following, we call $E_{j}$ the value of $V_{eff}$ in the Lagrange point $L_{j}$. And we use the names $E_{6}$ and $E_{7}$ for the values of $V_{eff}$ in the points $P_{1}$ and $P_{2}$ respectively. The numerical values of $E_{j}$ for our choice of the masses are: $E_{1}=-0.06511$..., $E_{2}=-0.05481$..., $E_{3}=-0.05921$..., $E_{4}=E_{5}=-0.04647$..., $E_{6}=-0.2714$... and $E_{7}=-0.5714$....
Figure 1: The effective potential in the horizontal plane. The green, blue and red curves are the equipotential lines to the energies $E_{1}$, $E_{2}$ and $E_{3}$, respectively. Some additional equipotential curves are drawn as grey lines. The black dots mark the extremal points of the potential.
In Fig.1, the equipotential curves to the energies $E_{1}$, $E_{2}$ and $E_{3}$ are given by the green, blue and red curves, respectively. The equipotential curves for some further energy values are plotted as grey curves. The extremal points of the potential are marked by black dots and labelled by their names given in the last paragraph.
Finally, we give the motivation for our choice of the mass ratio between the two galaxies. As already mentioned, we want to keep $M_{t}$ fixed at the value 0.2. And then, it is clear that we have chosen $M_{1}=M_{2}=0.1$ for the symmetric case in [21]. And it is also obvious that in the symmetric case we had $E_{2}=E_{3}$ and also $E_{6}=E_{7}$, see Fig.3 in [21]. In the present article, we want to have the system strongly asymmetric and all three $E_{j}$ values of the collinear Lagrange points as distinct as possible. So we have chosen the values of the masses such that the distance between $E_{1}$ and $E_{3}$ is approximately equal to the distance between $E_{3}$ and $E_{2}$. Now the reader might ask: Why not use $M_{2}$ close to 0.2 and correspondingly $M_{1}=M_{t}-M_{2}$ close to zero? Then the ratio between $M_{2}$ and $M_{1}$ would be enormous. The problem is that in such a case, $E_{3}$ and $E_{1}$ would be very close together, and the saddle point $L_{3}$ would have little importance for the dynamics.
3 Numerical results for the development and decay scenario of the NHIMs
3.1 Development scenario of the most important periodic orbits
From now on, we use the conserved numerical value $E$ of the Hamiltonian as the perturbation parameter of the system and study the development scenario of the dynamics as a function of $E$. Out of the index-1 saddles and out of the relative minima of $V_{eff}$ grow the most important periodic orbits of the system, which are partly also the most important periodic orbits within the NHIMs. Therefore, a detailed knowledge of the development scenario of these periodic orbits helps a lot to understand the development scenario of the NHIMs and, in the end, the development scenario of the whole dynamics. The important properties of these periodic orbits are displayed in the bifurcation diagram plotted in Fig.2 in an energy interval which is relevant for the topic of the present article.
Figure 2: Bifurcation diagram of the most important periodic orbits. Red curves represent horizontal orbits. Dark green curves represent vertical orbits, blue curves represent tangentially stable tilted loop orbits, and the magenta curve represents tangentially unstable tilted loop orbits. Pitchfork bifurcations are marked as open black squares, and the loss of normal hyperbolicity is marked by eight-pointed black stars.
The horizontal axis of the plot gives the $x$ coordinate of the periodic orbit at the moment when it crosses the plane $y=0$ of the position space in negative orientation. The vertical axis of the plot gives the value $E$ of the generating function $H$.
Out of each minimum of the potential grows three fundamental periodic orbits, two of them are horizontal, i.e. completely contained in the plane $z=0$ of the position space, and one of them goes mainly into the vertical direction. One horizontal orbit has a negative sense of rotation, and the other one has a positive orientation of rotation. The horizontal orbits with positive orientation are not important for the development scenario of the dynamics and are not discussed in the following, nor included in Fig.2 in order not to overload the plot. In contrast, all the periodic horizontal orbits with a negative orientation and with period one with respect to the negative intersection with the plane $y=0$ are central for the development scenario and are included in Fig.2 as the red lines. There is one red curve growing out of each of the potential minima. Note that each one of these two red curves runs through a symmetry-broken pitchfork bifurcation, whereby each one turns into 3 red curves for higher values of the energy. For the curve coming out of the minimum $P_{1}$, this happens for $x\approx-3,E\approx-0.063$, and for the one growing out of the minimum $P_{2}$, it happens for $x\approx 2.8,E\approx-0.074$. For the event of symmetry-broken pitchfork bifurcations, see section 20.3 and Fig. 20.3.2 in [26].
The approximately vertical periodic orbits are represented in Fig.2 as the dark green curves. Again, the $x$ value of the orbit at the moment of the negative intersection with the plane $y=0$ is plotted as a function of $E$.
Out of each one of the 3 collinear Lagrange points grows one horizontal Lyapunov orbit with negative orientation of rotation in the xyx-y plane, again represented in Fig.2 as red curves, and one approximately vertical Lyapunov orbit, again represented in the plot as dark green curves. In the following, we call lhj and lvj the horizontal Lyapunov orbit and the vertical Lyapunov orbit, respectively, coming out of the Lagrange point $L_{j}$. These Lyapunov orbits are the most important and central periodic orbits within the NHIMs created over the collinear Lagrange points. And these Lyapunov orbits are born at $E_{j}$ as normally hyperbolic and tangentially elliptic, where the words normally and tangentially refer to the NHIM surface.
Pitchfork bifurcations play an important role in the scenario, and they are indicated by open black squares in the plot. In such a pitchfork bifurcation, each one of the horizontal Lyapunov orbits splits off a pair of tilted loop orbits, which go out of the horizontal plane. The two members of such a pair are transformed into each other by a reflection in $z$ and are represented by a single blue curve in the Fig.2. For higher energy, these tilted loop orbits disappear in saddle-centre bifurcations with pairs of further tilted loop orbits (also represented by blue curves) being created in pitchfork bifurcations of the horizontal orbits coming out of the potential minima. The tilted loop orbits of this type created over the Lagrange point $L_{j}$ will be called ltj in the following. The orbit lh1 runs through a further pitchfork bifurcation at $E\approx-0.045$, where it splits off a further pair of tilted loop orbits which are represented by the magenta curve in the plot. These tilted loop orbits will be called lu1. With increasing energy, these new tilted loop orbits lu1 turn vertical very rapidly and are absorbed by the vertical Lyapunov orbit lv1 in an inverse pitchfork bifurcation at $E\approx-0.043$.
The tangential stability properties of the periodic orbits contained within the NHIMs and their changes during all these bifurcations can be visualised best in the restricted Poincaré maps of the NHIMs. They will be explained in detail below in the subsections describing the development scenarios of the NHIMs.
At still higher energy, the horizontal Lyapunov orbits coming out of the collinear Lagrange points are destroyed in saddle-centre bifurcations together with horizontal periodic orbits coming out of the potential minima or the horizontal orbits created in the broken pitchfork bifurcations suffered by the horizontal orbits coming out of the potential minima. Shortly before their destruction, the horizontal Lyapunov orbits lose their normal hyperbolcity; their loss of normal hyperbolicity is marked in the plot by a black eight-pointed star.
Now let us compare the actual Fig.2 with the Fig.11 in [21] for the corresponding symmetric case. Because of the symmetry in the old plot, we omitted the orbits coming out of $L_{3}$ and a large part of the orbits coming out of $P_{1}$. On the other hand, in the old plot, we included a few additional orbits which have turned out to be of little interest for the development scenario. In the new plot, these orbits are not included in order not to overload the plot. Therefore, to compare with the new asymmetric plot, we should imagine the old plot of the symmetric case augmented by the orbits coming out of $L_{3}$ and of $P_{1}$ and should delete the irrelevant orbits. Then we see a strong qualitative analogy between the two plots. There is only one essential difference which we point out now. In the new plot in Fig.2, we see two rather sharp corners in the red curves, one at $x\approx 0.3,E\approx-0.039$ and the other one at $x\approx 1.5,E\approx-0.035$. When the two masses become more similar and approach each other, then these two corners also approach each other in the middle and for equal masses, they touch and form a pitchfork bifurcation where lh1 absorbs a pair of other horizontal orbits coming from the potential minima. What we see in the asymmetric case in Fig.2 is a symmetry-broken pitchfork instead.
3.2 Development scenario of the internal dynamics of the NHIM1
(a) E=0.064E=-0.064
(b) E=0.046E=-0.046
(c) E=0.044E=-0.044
(d) E=0.042E=-0.042
(e) E=0.041E=-0.041
(f) E=0.04E=-0.04
Figure 3: Restricted Poincaré map of the NHIM1 for different energies
We show the internal dynamics of NHIMs by the Poincaré map restricted to the NHIM surface with the intersection condition $z=0$ where the intersection orientation is irrelevant because of the symmetry under $z$ reflection. In the full 4-dimensional Poincaré map, the NHIM is a 2-dimensional curved surface embedded in the 4-dimensional domain of the map having coordinates $x,y,p_{x},p_{y}$. To display the restricted map graphically, we show the projection of the NHIM surface into some appropriate 2-dimensional coordinate surface. Here, the choice of the $y-p_{y}$ coordinate plane gave the best graphical results. Therefore, we will use this projection for all restricted Poincaré maps throughout the article.
The central and most important NHIM is the one over the Lagrange point $L_{1}$. Therefore, we start with the graphical presentation of the development scenario of this NHIM. It is created at the energy $E_{1}$. Very close to this energy, the internal dynamics of the NHIM is a two degrees of freedom (2-dof) oscillator, and accordingly, the Poincaré plot in Fig.3(a) for E=0.064E=-0.064 looks like that the one of a 2-dof oscillator. The central fixed point (a green dot in the figure) represents the vertical Lyapunov orbit lv1, and the boundary (shown as a solid red curve) is the horizontal Lyapunov orbit lh1, which lies completely in the horizontal plane. For such low energies, both of these Lyapunov orbits are tangentially elliptic. Now we increase the energy and show plots of the qualitatively changed NHIM1 appearance after each important bifurcation event, which is related to an important bifurcation event of the Lyapunov orbits, presented in the previous subsection.
At an energy $E\approx-0.0638$, lh1 runs through its first pitchfork bifurcation, where it becomes tangentially hyperbolic and splits off a pair of tangentially elliptic tilted loop orbits. In the restricted map, these tilted loop orbits are represented by new fixed points which, at the bifurcation energy, come out of the boundary, i.e. out of lh1. In the figures, these tilted loop fixed points are represented by blue dots. The tilted loop orbits are surrounded by KAM islands, which grow larger with increasing energy. Fig.3(b) shows the situation for E=0.046E=-0.046. Here, the central island around lv1 is already rather small; the large majority of the area of the NHIM1 is occupied by the islands around the tilted loop orbits. In its first pitchfork bifurcation, lh1 has become tangentially unstable and has become part of a separatrix structure which divides the various islands. Strictly speaking, this separatrix structure is a fine chaos strip. However, it is so fine that in the figure it looks like lines. The tangential instability of lh1 is here still smaller than its normal instability. Accordingly, lh1 still belongs to the NHIM1.
At an energy a little below E=0.045E=-0.045, lh1 runs through its second pitchfork bifurcation, where it splits off a new pair of tangentially hyperbolic tilted loop orbits, while lh1 itself returns to tangentially elliptic. The Fig.3(c) for the energy E=0.044E=-0.044 shows how these new tilted loop orbits run along the line $y=0$ closer to the central point for increasing energy. The 2 fixed points in the map, which correspond to these 2 new tilted loop orbits, are marked by magenta dots. The unstable manifolds of the new tangentially unstable tilted loop orbits create a fine chaos strip, which in the figure looks like a separatrix structure. This separatrix divides all the various disjoint regular islands. First, we have the very small remnant of the central island around lv1. Then we have the island pair around the tangentially elliptic tilted loop orbits. And last, we have a new regular region of KAM curves close to the boundary (which is still lh1), which has now become tangentially elliptic again.
At the energy $E\approx-0.043$, the tangentially hyperbolic tilted loop orbits reach the central point and are absorbed by this central point in an inverse pitchfork bifurcation. Thereby, the central point turns tangentially hyperbolic and becomes the central point of the separatrix structure. The island around lv1 has been lost completely in this bifurcation. The Fig.3(d) shows the NHIM1 for E=0.042E=-0.042. We see how the separatrix structure has become a chaos strip of a size which is already clearly visible in the plot. In addition, we notice that the tilted loop islands start to disintegrate from the outside, i.e. from the neighbourhood of the chaos strip. Now we are in the energy region, which is most important for the whole development scenario. Therefore we show in Fig.3(e) the NHIM1 for the nearby energy E=0.041E=-0.041, even though no important bifurcation event happens during this further increase of the energy. Now the separatrix has already turned into a large-scale chaos region. It is important to emphasise that up to this energy, all elements in NHIM1 are still normally hyperbolic, i.e. in all points of the NHIM1, the normal instability is still larger than the tangential instability. This also holds for the large chaos region.
At $E\approx-0.04035$, a very drastic event happens to the NHIM1. Here, the Lyapunov orbit lh1 loses its normal hyperbolicity. Accordingly, for higher values of the energy, lh1 no longer belongs to NHIM1, and it can no longer act as the boundary of NHIM1. At this energy, NHIM1 starts to decay. At first sight, the plot of NHIM1 for E=0.04E=-0.04 in Fig.3(f) does not look very different from the plot in Fig.3(e). Nevertheless, there are some important qualitative differences which we have to point out. As already mentioned, lh1 is no longer the boundary of NHIM1. Therefore, this orbit has been included in the plot as a broken line only. Now the large chaos strip no longer has any boundary curve on its outside, which prevents trajectories from diffusing chaotically to the outside in the long run and leaving the NHIM1 tangentially. Therefore, strictly speaking, the chaos strip is no longer invariant and accordingly no longer a part of the NHIM1. Trajectories only stay temporarily in the apparent chaos strip of the NHIM1. We can say that this apparent chaos strip only has a transient existence as an invariant surface. For more details on this type of transient behaviour on decaying NHIMs, see [23].
3.3 Development scenario of the NHIM3
Next, let us have a look at NHIM3 over the Lagrange point $L_{3}$. This NHIM3 is created at the energy $E_{3}\approx-0.05921$. For energies a little larger than $E_{3}$, also this NHIM has an internal dynamics of a 2-dof oscillator, and accordingly, the reduced Poincaré map shows the corresponding pattern of concentric ellipses, where the central fixed point (again represented by a green dot in the figures) represents the vertical Lyapunov orbit lv3 and the boundary (again represented by a red solid curve in the figure) is the horizontal Lyapunov orbit lh3. It is not worth showing a plot for a very low energy.
(a) E=0.045E=-0.045
(b) E=0.043E=-0.043
(c) E=0.042E=-0.042
(d) E=0.041E=-0.041
(e) E=0.04E=-0.04
Figure 4: Restricted Poincaré map of the NHIM3 for different energies.
At an energy $E\approx-0.055$, lh3 runs through a pitchfork bifurcation where it becomes tangentially hyperbolic and splits off a pair of tangentially elliptic tilted loop orbits, which in the reduced Poincaré map are the centres of islands and are again represented by blue dots in the figures. The Fig.4(a) shows the case of E=0.045E=-0.045 where the islands around the tilted loop orbits have already grown so large that they occupy approximately half of the area of NHIM3. Unlike lh1 and as we know already from subsection 3.1 and Fig.2, lh3 does not run through a further pitchfork bifurcation and remains tangentially hyperbolic as long as it belongs to NHIM3. Note that for the present energy, lh3 has a higher normal instability than its tangential instability; therefore, it is still normally hyperbolic and forms part of NHIM3. It is the outer part of a separatrix structure which it forms together with the separation line between the central island and the tilted loop islands. Strictly speaking, this separatrix structure is again a fine chaos strip; however, so fine that it appears as lines under the resolution of the graphics.
We are already in the interesting energy region, therefore we increase $E$ in small steps and present in Fig.4(b) the reduced map of NHIM3 for E=0.043E=-0.043 even though the qualitative structure remains the same. We see that in comparison with Fig.4(a) many invariant lines from Fig.4(a) are broken into corresponding very fine secondary structures in Fig.4(b). This is a strong indication that we are already close to large-scale chaos and decay of the NHIM3.
In Fig.4(c), we proceed to E=0.042E=-0.042 and see a very drastic change in comparison to Fig.4(b). As we already know from Fig.2, for energies a little smaller than -0.042, several important bifurcations of the important periodic orbits in NHIM3 occur. First, at $E\approx-0.0424$ lh3 loses its normal hyperbolicity and can no longer be a part of NHIM3. Here, the tangential instability has become larger than the normal instability. Second, at an energy extremely close to this event, at $E\approx-0.04239$, lh3 is destroyed in a saddle-centre bifurcation when it collides with another horizontal periodic orbit grown out of the potential minimum $P_{1}$. Accordingly, the outer boundary and the separatrix structure of the NHIM3 are lost. Third, at an energy $E\approx-0.04232$, the tilted loop orbits from NHIM3 collide with another pair of tilted loop orbits grown out of $P_{1}$ (see again Fig.2). Thereby, the tilted loop orbits together with their surrounding islands are lost from NHIM3. The only part of NHIM3 which persist this bifurcation process is the inner part of the central island around the vertical Lyapunov orbit lv3.
In the region where the tilted loop islands have been destroyed, we find transient remnants of the NHIM3 in analogy to what we saw before in Fig.3(f) for NHIM1. In Fig.4(c), one transient trajectory in this decaying region is included by the orange crosses. We can imagine that in the decay of the 2-dimensional NHIM surface, lower-dimensional invariant subsets remain, and this collection of lower-dimensional invariant sets can hold general trajectories in their neighbourhood at least temporarily. Of course, these lower-dimensional unstable invariant sets also have stable and unstable manifolds of corresponding lower dimension, which can hold very particular incoming trajectories forever. But our numerical methods adapted to search for 2-dimensional NHIM surfaces and their 3-dimensional stable manifolds do not catch these lower-dimensional exceptional subsets. We only notice that general trajectories suffer some time delay in this region of the domain of the map.
In Fig.4(d), we display the NHIM3 for E=0.041E=-0.041. With increasing energy, the size of the central island shrinks slowly. Here in the outer transient region, we encountered something instructive. We started a trajectory on the decaying NHIM surface at the initial coordinates $y=0$ and $p_{y}=-0.55$. This point happens to lie very close to the tangentially unstable central periodic point of a secondary 5:11 resonance. First, the iterates lie close to this periodic orbit before they deviate strongly for higher iterations and diffuse away. In total, we could follow this transient trajectory on the remnants of the NHIM3 for 125 iterations. This transient trajectory is again plotted by orange crosses. In Fig.4(e), we proceed to E=0.04E=-0.04. Compared to the previous part, there is little difference. Most interesting in this figure is the outer transient part. It is a single trajectory starting at $y=0$ and $p_{y}=-0.538$. The algorithm could find 81 iterations, the last one is the point near $y=4$, $p_{y}=-0.107$. We could not find a next iterate inside of the region belonging to the logical continuation of the NHIM3 to higher energies.
For still higher energies, the central island of the NHIM3 shrinks slowly but remains. For $E\approx-0.028$ lv3 becomes tangentially unstable, but it remains normally hyperbolic even up to positive energies. The continuation of NHIM3 for such high energies is no longer relevant for the theme of the present article; therefore, we do not follow it any longer.
3.4 Development scenario of the NHIM2
Figure 5: Restricted Poincaré map of the NHIM2 at E=0.035E=-0.035.
Finally, let us turn to NHIM2. It is created at $E_{2}\approx-0.05481$, and also this NHIM starts with an internal dynamics of a 2-dof oscillator. Also, here the horizontal Lyapunov orbit lh3 runs through a pitchfork where it splits off a pair of tilted loop orbits, it happens at $E\approx-0.046$, see again Fig.2. With increasing energy, the islands around the tilted loop orbits grow very slowly, and we must reach rather high energies to obtain some interesting plot of the restricted map of NHIM2. In Fig.5, we show the plot for E=0.035E=-0.035. Of course, since its pitchfork bifurcation, lh2 is unstable and together with the division between the central island and the tilted loop islands, it forms a separatrix structure, which, strictly speaking, is a very fine chaos strip. The apparent intersections between the various curves in the tilted loop islands are a projection effect. Here, the projection of the curved NHIM2 surface into the $y-p_{y}$ plane is no longer 1:1; however, we always want to use the same projection for all plots of the restricted map.
At $E\approx-0.032$ lh2 first loses its normal hyperbolicity and immediately afterwards is destroyed in a saddle-centre bifurcation when it collides with another horizontal periodic orbit growing out of the potential minimum $P_{2}$. At $E\approx-0.02997$, the tilted loop orbits in NHIM2 first lose normal hyperbolicity and immediately afterwards are destroyed in saddle-centre bifurcations when they collide with other tilted loop orbits grown out of $P_{2}$, see again Fig.2. The vertical orbit lv2 remains normally hyperbolic up to very high energy, and accordingly, also a remnant of the NHIM2 around lv2 remains. Qualitatively, the scenario of NHIM2 is a mirror image of the scenario of NHIM3, even though the analogous individual events happen at higher values of $E$. So we see in which sense a qualitative mirror symmetry of the system is conserved, also when the symmetry between the masses is broken.
3.5 Some comments on the NHIM development scenarios
Important events happen in NHIM1 and also in NHIM3 in the same relatively small energy interval between E=0.043E=-0.043 and E=0.04E=-0.04. With increasing energy in NHIM3, the loss of normal hyperbolicity happens before the development of large-scale chaos in the internal dynamics. We see large-scale chaos in the surroundings of the remnant parts of NHIM3 only as a transient effect. In contrast, in the internal dynamics of NHIM1, we observe the development of large-scale chaos before the beginning of any decay of this NHIM. The difference is that NHIM3 starts its decay because the normal instability is becoming small, whereas in NHIM1, the development of the growing tangential instability is important for the beginning of the decay.
When we compare the development scenario of NHIM1 on one hand with the development scenarios of NHIM2 and NHIM3 on the other hand, then we notice the following qualitative difference. On NHIM1, the horizontal Lyapunov orbit runs through two consecutive pitchfork bifurcations, where in the second one, a pair of unstable tilted loop orbits is created, while the horizontal Lyapunov orbit returns to tangential stability. This second pair of tilted loop orbits turns vertical and is finally absorbed by the vertical Lyapunov orbit in an inverse pitchfork bifurcation. Thereby, the vertical Lyapunov orbit becomes tangentially unstable, and this event triggers the creation of large-scale chaos inside of NHIM1.
In contrast, in NHIM2 and in NHIM3, tangentially unstable tilted loop orbits are never created, which has two consequences. First, the horizontal Lyapunov orbit remains tangentially unstable, which favours the beginning of the decay of these NHIMs from the separatrix structure, which contains lh2 or lh3, respectively. Second, the vertical Lyapunov orbit remains tangentially stable, which favours the survival of the central island of these NHIMs up to very high energy.
We made a surprising observation of the connection of these properties with the behaviour of the time periods of the Lyapunov orbits as a function of the energy. In Fig.6, we plot these time periods as a function of the energy. As before, the red curves belong to the horizontal Lyapunov orbits, and the green curves belong to the vertical Lyapunov orbits. The blue curves belong to the tangentially stable tilted loop orbits. And the magenta curve belongs to the tangentially unstable tilted loop orbits. The solid curves belong to NHIM1, the dotted curves belong to NHIM2 and the broken curves belong to NHIM3. In NHIM1, the periods of the horizontal and the vertical Lyapunov orbit cross each other; one pitchfork bifurcation happens energetically before the crossing, and the other one happens after the crossing. In NHIM2 and NHIM3, we do not find any analogous crossing of the horizontal and vertical periods.
Figure 6: Periods of Lyapunov orbits and tilted loop orbits as a function of the energy.
Now the reader might ask whether this observation is a strange feature of the particular example or whether it is common. Interestingly, the experience with other systems suggests that this behaviour is common. In the system of a test particle in a barred galaxy [27] and for the dynamics of a test particle in a double-barred galaxy [22], we found two consecutive pitchfork bifurcations creating first a pair of tangentially stable tilted loop orbits and later a pair of tangentially unstable tilted loop orbits. Also, a crossing of the periods of the horizontal and the vertical Lyapunov orbits occurs between these two pitchfork bifurcations. However, in these two examples, the tilted loop orbits come out of the vertical Lyapunov orbit. On the other hand, in the system of a test particle in the gravitational field of an open star cluster [28], there is no crossing of periods, and there is only a single pitchfork bifurcation creating a pair of tangentially stable tilted loop orbits.
In a classical model for a hydrogen atom in a rotating external field [29], we find the following scenario. First the horizontal Lyapunov orbit splits off a pair of tangentially stable tilted loop orbits. Later, we have a crossing of the periods of the horizontal and vertical Lyapunov orbits. However, at an energy a little higher, the horizontal Lyapunov orbit collides with the singularity of the Coulomb system at the origin, and this event throws the scenario out of its usual path. At an energy still a little higher, the horizontal Lyapunov orbit is absorbed by an orbit coming out of the central potential hole in an inverse period doubling bifurcation. So the horizontal Lyapunov orbit ends its existence before it has the opportunity for a further pitchfork bifurcation.
The conclusion of this subsection is the following. We have seen the creation of tangentially unstable tilted loop orbits only in cases where the periods of the horizontal and the vertical Lyapunov orbits cross. The system of the two dwarf galaxies is a nice example where we see in the same system the case with crossing and the case without crossing. Certainly, the crossing of the periods indicates some kind of 1:1 resonance between the horizontal and the vertical motion. And such a resonance is needed for the creation of tilted loop orbits out of the Lyapunov orbits. Unfortunately, so far we are not yet able to give a detailed explanation of all these connections.
4 Heteroclinic connections and the appearance of large-scale chaos
Now we relate the NHIM development scenario seen in section 3 to some particularly interesting classes of trajectories and their properties, namely, to heteroclinic trajectories between the 3 NHIMs. We do it for the energy values 0.043-0.043, 0.042-0.042 and 0.04-0.04, i.e. exactly for the energy range where the most interesting events of the whole development scenario happen. Because of the easier pictorial presentation, we begin with horizontal trajectories. The only horizontal trajectories within the NHIMs are the respective horizontal Lyapunov orbits, which have already served as boundaries of the NHIM representations in the plots of the restricted map. Fig.7 shows the three horizontal Lyapunov orbits lh1, lh2 and lh3 (plotted as red lines) in the horizontal plane of the position space for E=0.043E=-0.043. These three orbits are identified in the plot by their order along the $x$ axis. We know already from Fig.2 that this energy value is very close to the highest energy where all the important fundamental periodic orbits still exist. The figure contains, in addition, the projection into the horizontal plane of the vertical Lyapunov orbits (dark green curves) and of the tilted loop orbits (blue curves). The orbit lu1 disappears already exactly at this energy when it collides with lv1. Therefore, it looks very similar to lv1 at this energy, and it is not included in the figure because it would be extremely difficult to distinguish lu1 and lv1. Note that lh3 and lh1 come very close to each other, while lh2 keeps a greater distance from lh1. Note that the two tilted loop orbits of each NHIM have identical projections into the horizontal plane.
Figure 7: Projection of the simplest and most important periodic orbits at E=0.043E=-0.043 into the horizontal plane of the position space. The red curves are horizontal Lyapunov orbits, the green curves are vertical Lyapunov orbits , and the blue curves are tangentially stable tilted loop orbits.
The trace of any one of these periodic orbits is symmetric under $y$ reflection; the $y$ reflection only reverses the time orientation of the orbits. This is a consequence of the symmetry from Eq.6 under which the whole dynamics is invariant.
4.1 Heteroclinic trajectories
Figure 8: Heteroclinic trajectories between various substructures of the NHIMs at E=0.043E=-0.043. (a) : A horizontal heteroclinic trajectory going from lh3 to lh1. (b) : A horizontal heteroclinic trajectory going from lh3 to lh2. (c) : A horizontal heteroclinic trajectory going from lh2 to lh1. (d) : Projection into the horizontal plane of a heteroclinic trajectory going from lv3 to the outer part of the tilted loop island of NHIM1. (e) : Projection into the horizontal plane of a heteroclinic trajectory going from lt3 to the outer part of the central island of NHIM2.
First, we have a look at heteroclinic orbits between the horizontal Lyapunov orbits. They are the only horizontal heteroclinic orbits between the various NHIMs. In Fig.8(a), we show a typical example of a heteroclinic trajectory which starts on lh3 and ends on lh1. In Fig.8(b), we show a heteroclinic trajectory going from lh3 to lh2 and Fig.8(c) presents a heteroclinic trajectory going from lh2 to lh1. Of course, there is an infinity of further horizontal heteroclinic trajectories between the various horizontal Lyapunov orbits performing all kinds of loops in the regions of the NHIMs before settling down to one particular Lyapunov orbit in the future and in the past.
We also show two examples of heteroclinic trajectories going out of the horizontal plane. However, in the figures, we plot their projections into the horizontal plane. Fig.8(d) presents a trajectory starting on lv3 and ending on the outer part of the tilted loop island of NHIM1 close to the separatrix. Fig. 8(e) presents a trajectory starting on lt3 and ending on the outer part of the central island of NHIM2.
Now we increase the energy a little to -0.042. We know already that for this energy, the periodic orbit lh3 has already disappeared in a saddle-centre bifurcation included in Fig.2. Accordingly, there is no longer any horizontal trajectory inside the remaining part of NHIM3 and therefore also no horizontal heteroclinic trajectory starting or ending on NHIM3. However, we have the transient chaotic region presented in the NHIM3 plot of Fig.4(c). And some of these transient trajectories around the remnants of NHIM3 have a projection into the horizontal plane which qualitatively looks similar to lh3 at smaller energies, i.e. it looks as if there would be a logical continuation of lh3 to higher energies. From the neighbourhood of such transient trajectories, heteroclinic trajectories to other NHIMs can start. Fig.9(a) shows the projection into the horizontal plane of the position space of a typical trajectory going from the transient part of NHIM3 to NHIM2. Its initial point lies very close to the 26th point of the transient trajectory shown in Fig.4(c), its coordinates in this figure are $y\approx 0.57,p_{y}\approx-0.39$. On the NHIM2, this heteroclinic trajectory approaches the outer part of the central island in NHIM2.
Fig.9(b) shows the projection into the horizontal plane of the position space of a trajectory going from the transient part of NHIM3 to NHIM1. The initial point of this trajectory lies extremely close to the initial point of the trajectory shown in Fig.9(a). Compare in the figures the initial point and the first loop of these two trajectories in the region of NHIM3. However, in the long run, these two trajectories are completely different. This demonstrates the extreme dependence on the initial conditions of trajectories starting in the chaotic region, a typical indication of chaos. This heteroclinic trajectory approaches the NHIM1 in the region between the moderate-sized chaos strip created around the stable and unstable manifolds of the central hyperbolic fixed point corresponding to the vertical Lyapunov orbit lv1 and the outermost layer of the tilted loop islands, which is already broken into secondary structures and chaos strips. It establishes a connection between the transient chaotic and decaying part of NHIM3 and the part of NHIM1 which is just in the process of turning chaotic at this stage of the development scenario.
Fig.9(c) shows a trajectory starting in the chaos region of NHIM1 and coming close to the outer transient parts of NHIM3. Fig.9(d) shows a trajectory starting close to lv1, i.e. in the centre of the chaotic region of NHIM1 and approaching NHIM2.
Figure 9: Projection into the horizontal plane of four heteroclinic trajectories between the NHIMs at E=0.042E=-0.042. Part (a): A heteroclinic trajectory from the transient part of NHIM3 to NHIM2. part (b): A heteroclinic trajectory from the transient part of NHIM3 to NHIM1. part (c): A heteroclinic trajectory from the large-scale chaotic region of NHIM1 to NHIM3. part (d): A heteroclinic trajectory from lv1 to NHIM2.
Next, we go to E=0.04E=-0.04 where NHIM3 and NHIM1 have both started with their decay. The Fig.10(a) shows a heteroclinic trajectory starting on the transient part of NHIM3 and going to NHIM1. And the Fig.10(b) shows a heteroclinic trajectory starting on the transient part of NHIM1 and ending on NHIM3.
Figure 10: Projection into the horizontal plane of two trajectories at E=0.04E=-0.04. Part (a): A heteroclinic trajectory going from the transient part of NHIM3 to NHIM1. Part (b): A homoclinic trajectory going from the transient part of NHIM1 to NHIM3.
4.2 The horizontal horseshoe
The properties of homoclinic and heteroclinic connections can be illustrated best by the construction of horseshoes in some appropriate Poincaré map. We have already seen in the previous subsection that horizontal heteroclinic trajectories help a lot to illustrate heteroclinic connections between the various NHIMs. In addition, we can only plot Poincaré maps restricted to some invariant 2-dof subsystem. Therefore, we construct now a Poincaré map for the horizontal subsystem and illustrate the creation of heteroclinic connections between the NHIMs in this 2-dimensional map. For the construction of this Poincaré map, we use the intersection condition $y=0$ with negative orientation. Remember that it is the same intersection condition which we have used before in the construction of Fig.2. Therefore, we can relate important structures in the horizontal Poincaré map with particular periodic orbits contained in Fig.2 by just comparing the $x$ coordinates. The domain of the horizontal Poincaré map is the $x-p_{x}$ plane.
As already mentioned before, the only horizontal trajectories within the NHIMs over the collinear Lagrange points are the respective horizontal Lyapunov orbits. And in the horizontal Poincaré map, they are represented by fixed points on the line $p_{x}=0$. Their $x$ coordinate can be read off Fig.2. We use the energy value -0.043 because this is the highest energy where the three horizontal Lyapunov orbits lh1, lh2, and lh3 still exist, are still normally hyperbolic and still act as boundary curves of the NHIMs, as explained in section 3. Therefore, it is clear that the corresponding fixed points in the horizontal Poincaré map are hyperbolic fixed points. In the Poincaré plot, these three fundamental fixed points are marked by the black open circles and are called lh1, lh2 and lh3 respectively, exactly equal to the Lyapunov orbits which they represent.
We calculate initial segments of the stable and unstable manifolds $W^{s}(lhj)$ and $W^{u}(lhj)$ of the three fundamental fixed points and plot them into the $x-p_{x}$ plane, i.e. into the domain of the horizontal map. The result is Fig.11. From the trajectory plots in the horizontal position space shown in the previous subsection, it is obvious that we have frequent violations of the transversality of the intersection condition $y=0$. Therefore, the stable and unstable manifolds are torn into many disjoint fragments. In the plot $W^{s}(lh1)$, $W^{u}(lh1)$, $W^{s}(lh2)$, $W^{u}(lh2)$, $W^{s}(lh3)$ and $W^{u}(lh3)$ are plotted in the colours orange, green, red, blue, magenta and cyan, respectively. The black curves are the boundary of the energetically accessible region of the $x-p_{x}$ plane at E=0.043E=-0.043.
Figure 11: Horseshoe at E=0.043E=-0.043 in the invariant horizontal subsystem built up by the stable and unstable manifolds of the horizontal Lyapunov orbits sitting over the collinear Lagrange points. In the plot $W^{s}(lh1)$, $W^{u}(lh1)$, $W^{s}(lh2)$, $W^{u}(lh2)$, $W^{s}(lh3)$, and $W^{u}(lh3)$ are plotted in the colours orange, green, red, blue, magenta, and cyan, respectively. The fixed points representing the Lyapunov orbits are marked by black open circles. The black curves are the boundary of the energetically accessible region at E=0.043E=-0.043. Some KAM islands are included as grey curves.
Also included in grey colour are some KAM islands around tangentially stable periodic orbits. With the help of Fig.2, the stable elliptic fixed points in the centres of these islands located at $x\approx-2.7$, $x\approx 3.4$ and $x\approx 5.8$ respectively can be identified with three horizontal periodic orbits growing out of the potential minima.
In the figure, we make an interesting observation. Between the region around the Lagrange points $L_{1}$ and $L_{2}$, there exists an amazingly large KAM island which reaches almost to the energetic boundary of the domain of the map. This makes heteroclinic connections to the NHIM2 relatively difficult. In comparison, the KAM island between the regions around the Lagrange points $L_{1}$ and $L_{3}$ is a lot smaller and leaves more room for heteroclinic connections to NHIM3. This observation makes it understandable that NHIM1 has more coordination of its development scenario with NHIM3 than with NHIM2. Of course, for very high iterations of the initial segments of the stable and unstable manifolds, the ones from any one of the NHIMs accumulate against the ones from all other stable or unstable manifolds respectively of all invariant subsets of the dynamics. However, the corresponding heteroclinic trajectories might need a very long time to come from the direct neighbourhood of the initial invariant subset to the neighbourhood of the final invariant subset. On their way, these trajectories perform a lot of complicated intermediate loops.
It seems that for any efficient coordination of the development scenarios of NHIMs, we need heteroclinic trajectories which proceed along simple and fast trajectories between the direct neighbourhoods of these NHIMs.
4.3 Some remarks on transients and heteroclinic connections
The existence of the heteroclinic trajectories starting or ending on transient parts of decaying NHIMs might lead us to the following interesting speculation: The beginning of the decay of NHIM3 and the formation of the corresponding transient chaotic regions is synchronised with the formation of a large-scale chaotic region on NHIM1. Such a synchronisation is supported by the existence of heteroclinic trajectories like the one presented in Fig.9(b).
After having seen numerical evidence for transient effects and their influence on the general dynamics, let us finish this section with some further remarks on the decay of NHIMs and the related appearance of transient behaviour in the region of the former NHIM. As long as the NHIM is a smooth 3-dimensional invariant surface of the flow in the 5-dimensional energy shell (or a corresponding 2-dimensional surface in the 4-dimensional domain of the map), its stable manifold is a smooth 4-dimensional surface in the energy shell (or a corresponding 3-dimensional surface in the map). It must have codimension 1 in the energy shell and also in the domain of the map. Then a 1-dimensional curve $C$ in an appropriate region of the energy shell close to the NHIM intersects the local branch of the stable manifold transversally. The curve $C$ can be chosen to lie in the domain of the map, and then it also serves in the domain of the map for the same purpose. This is the basic idea behind the search strategy for the NHIM as explained in all details in [20].
Now imagine that the NHIM decays and that in some region only a lower-dimensional fractal collection of remnants of the NHIM remains. These remnants certainly contain unstable periodic orbits. The stable manifolds of these remnants have a lower dimension, i.e. higher codimension than the stable manifold of the former complete NHIM. Then, in general, the curve $C$ will no longer intersect exactly the stable manifolds of the remnants. Accordingly, the search strategy for the NHIM mentioned above will no longer find these lower-dimensional remnants. However, if the curve $C$ is chosen in an appropriate region, it still contains points leading to a high time delay in the neighbourhood of the remnants. This is the explanation of the transient effects. To find again the stable manifolds of such lower-dimensional remnants of NHIMs, it would be necessary to scan initial conditions not only along a 1-dimensional curve but on a higher-dimensional surface, where the dimension of this surface of initial conditions should be larger than or at least equal to the codimension of the stable manifold we want to catch.
5 Indicator functions
The phase space indicator functions are useful tools to analyse and get insights into the dynamics of systems with many degrees of freedom. A review of the topic and several examples can be found in [30, 31, 32]. In the present study, we are going to calculate the indicator based on the delay time. This quantity is an appropriate indicator function to study the phase space of open scattering systems [20, 24]. Some recent studies of the phase space and the transport using indicator functions for a similar model in 2 dimensions are in the reference [33].
To calculate the delay time indicator, we select the initial conditions on the canonical conjugate plane $xp_{x}$ with $y=0$, $z=0$, and $p_{z}=0$ for different values of the energy $E$, see the Fig.12. The white regions on the plots correspond to forbidden regions for the value of $E$. The lines with maximal values of the delay time indicator correspond to intersections of the initial conditions with the stable and unstable manifolds of the NHIMs, the yellow smooth regions are the intersections with the KAM stable tori, and the green regions correspond to initial conditions that escape fast to infinity. As the value of the $E$ increases, the structure of the homoclinic tangle changes. Let us notice that the initial conditions for the panel Fig.12(a) and Fig.11 are the same. We appreciate the match between the stable and unstable manifolds of the NHIMs and the KAM stable islands on both types of plots using different visualisation techniques.
a) E=0.043E=-0.043
b) E=0.042E=-0.042
c) E=0.041E=-0.041
d) E=0.04E=-0.04
e) E=0.039E=-0.039
f) E=0.037E=-0.037
Figure 12: Delay time for different values of $E$ and $z=0$
To continue with the analysis of the phase space of the system, we consider initial conditions on the canonical conjugate plane $xp_{x}$, $y=0$, $p_{z}=0$ and a fixed value of the energy E=0.0043E=-0.0043, but different values of initial $z$. The Fig.13 shows the intersections of the heteroclinic tangles of the NHIMs and the KAM tori with the set of initial conditions. The phase space has a similar structure for different values of $z=1,2,3$. The structure of the set of heteroclinic intersections and KAM islands looks very similar to the case with $z=0$ in Fig.12(a). However, we can see a change in the forbidden region for $z=4$; the set of initial conditions has two components.
(a) $z=1$
(b) $z=2$
(c) $z=3$
(d) $z=4$
Figure 13: Delay time indicator function for different values of $z$ and E=0.043E=-0.043
In the Fig.14, there is a fate map for the same initial conditions as Fig.12(a). For this plot, we only calculate the trajectories forward in time. The blue regions correspond to trajectories that escape to the interaction region with $x>0$, the green region to trajectories that stay trapped in the interaction region and the yellow region to trajectories that escape with $x<0$. We can see agreement betwee the delay time plot and the Poincaré map in Figs.12(a) and 11.
Figure 14: Fate map for E=0.043E=-0.043, $z=0$. The trajectories with yellow initial conditions escape to the left side, the dark blue ones escape to the right side, and the green ones stay trapped in the central region.
6 Conclusions and final remarks
The system analysed in the present article is an instructive example of Hamiltonian dynamics with 3 not symmetry-related NHIMs, and for the effects of coordination of the development scenario of the NHIMs. A new feature in the present article is the inclusion of tangentially transient behaviour on decaying NHIMs into the formation of the global scenario. We saw indications that also the decaying transient parts of NHIMs give important contributions to the heteroclinic tangle, i.e. to the global chaotic invariant set created by the NHIMs and their stable and unstable manifolds.
In the study of the development scenario of the NHIMs and the comparison with the corresponding left-right symmetric system investigated in [21], we made the following interesting observation: Also, when we destroy the symmetry between the two outer NHIMs by the transition to different masses of the two dwarf galaxies, there remains a partial qualitative symmetry in the development scenarios of the two outer NHIMs. This partial symmetry includes the sequence of the important bifurcations and their order for the Lyapunov orbits contained in the individual NHIMs. But it does not include the coordination of the creation of chaos. Here, in this example, only NHIM1 and NHIM3 make a coordinated development into the direction of chaos; NHIM2 does not participate in this coordination. This happens even though we have heteroclinic trajectories between any pair of NHIMs. The horizontal Poincaré map presented in subsection 5.2 gives at least some indication of the relatively weak heteroclinic coupling of NHIM2 to the other two fundamental NHIMs of the system.
It seems that NHIM1 and NHIM3 become susceptible to tangential instabilities at a similar energy interval, while NHIM2 is not yet ready for the development of large-scale chaos. Of course, in the symmetric case studied in [21], NHIM2 has exactly the same development scenario as NHIM3, and the two NHIMs develop large-scale chaos in exactly the same energy interval. The horizontal horseshoe plot of Fig.11 suggests that the large-scale KAM island over the potential minimum $P_{2}$ prevents a rapid and efficient heteroclinic connection between NHIM1 and NHIM2 along horizontal trajectories. In our collection of heteroclinic orbits, we do not have a single one which starts or ends on NHIM2, runs over the potential hole at $P_{2}$ and does not go to $z$ values far away from $z=0$. As we have seen in subsection 5.1, long and complicated trajectories can do it. As Figs.8(e) and 9(a) show, also trajectories going to the outside and circling around $L_{4}$ or $L_{5}$ can do it. But these long trajectories seem to be little efficient in the creation of coordination effects in the development scenarios between NHIMs. Of course, also trajectories using the vertical degree of freedom can establish heteroclinic connections, remember the trajectory from Fig.9(d). However, we have seen here (and also in other model systems) that in systems with a $z$ reflection symmetry and accordingly an invariant horizontal subsystem, the energy exchange between the vertical degree of freedom and the horizontal degrees of freedom is rather slow. This prevents an efficient coupling between the NHIMs via trajectories having a large part of their energy in the vertical motion. All these features prevent any efficient coupling of the development scenario of NHIM2 to the development scenario of the other two NHIMs.
One important aspect of the development scenario, which we can not yet explain precisely, but which certainly deserves a further more fundamental investigation, is the following consideration: For stable and unstable manifolds of NHIMs, there holds a foliation theorem which loosely speaking says that these manifolds are the Cartesian product of a line with the internal NHIM structure. Thereby, these manifolds transport the internal NHIM structure along them to all regions of the phase space heavily influenced by the homoclinic/heteroclinic tangle of the NHIMs. As a consequence, the whole homoclinic/heteroclinic tangle is at least qualitatively similar to a Cartesian product of a lower-dimensional homoclinic/heteroclinic tangle (in our case, the horizontal horseshoe presented in subsection 5.2) and the internal NHIM structures. This property has been noticed before and investigated in more detail in [34]. Also, the system investigated in this reference is 3-dof and has, in a natural form, an invariant horizontal 2-dof subsystem. Such a tendency of higher-dimensional homoclinic/heteroclinic tangles to show an approximate product structure has also been observed by another group in [35, 36, 37]. This tendency seems to be common and needs a more profound explanation.
The appearance of the tangential transient effects shows that, also after the beginning of its decay, the remnants of a NHIM have an effect on their environment in the phase space and on the formation of the homoclinic/heteroclinic tangle very similar to the effect the complete NHIM had before the start of the decay. The fragments and transient parts continue without heavy qualitative differences.
The NHIMs sitting over the Lagrange points of the effective potential direct the transport of test particles over these saddle points. In the interpretation of our model as describing two dwarf galaxies in interaction, this model describes the exchange of stars between the two galaxies (motion over the central Lagrange point $L_{1}$) or the loss of stars from the binary galaxy system (motion over the outer Lagrange points $L_{2}$ and $L_{3}$). The projection of the unstable manifolds of the NHIMs into the position space then gives the position of tidal arms and tidal bridges going out from the dwarf galaxies. Of course, our model system could also be interpreted as a simple model for two nearby globular clusters in interaction. Here, it is interesting to mention that in the astronomical literature, there is a claim of observational indications for a tidal bridge between the interacting globular clusters NGC5024 and NGC5053 [38].
Acknowledgment
We thank DGAPA-UNAM for financial support under grant number IG-101122 and CONACyT/SECIHTI for financial support under grant number 425854.